### v1

# DreamGarden: A Designer Assistant for Growing Games from a Single Prompt

Abstract.
Coding assistants are increasingly leveraged in game design, both generating code and making high-level plans. To what degree can these tools align with developer workflows, and what new modes of human-computer interaction can emerge from their use? We present DreamGarden, an AI system capable of assisting with the development of diverse game environments in Unreal Engine. At the core of our method is an LLM-driven planner, capable of breaking down a single, high-level prompt—a dream, memory, or imagined scenario provided by a human user—into a hierarchical action plan, which is then distributed across specialized submodules facilitating concrete implementation. This system is presented to the user as a garden of plans and actions, both growing independently and responding to user intervention via seed prompts, pruning, and feedback. Through a user study, we explore design implications of this system, charting courses for future work in semi-autonomous assistants and open-ended simulation design.
1. Introduction
A recent boom in generative AI has led to the widespread adoption of assistive AI tools by human creators both casual and professional in generating text, code, images and 3D assets. The AI models underpinning these tools are generally large neural networks, trained to predict or denoise next tokens in corpora of human-generated content scraped from the internet. Large language models are perhaps most ubiquitous, existing most commonly in the form of chatbots, predictive text models which have been fine-tuned via specialized datasets and human feedback to be generally “helpful” and excel in following user requests (OpenAI, 2024a; Touvron et al., 2023; Jiang et al., 2023). In the visual domain, diffusion models have brought on similarly impressive capabilities, allowing anyone to create visually convincing results from a single prompt (Rombach et al., 2022; Nichol et al., 2021; Ramesh et al., 2022). Most of these assistants respond however only to direct user prompting, and do not work autonomously toward a high-level goal.
In game design, procedural content generation (PCG) has long been leveraged to various ends; in turn effectively allowing for compression of large environments into generative algorithms to be executed at runtime (particularly in the early days of digital games, when disk space was a significant bottleneck), affording designers a new means of expression and a higher level of abstraction when designing content, accelerating the design process, and producing novel and surprising artifacts leading to increased replayability from the perspective of the player. PCG via machine learning is an active area of research, though industry professionals tend to still rely on more symbolic algorithms that provide finer-grained control and a greater degree of expressivity and interpretability.
In this work, we introduce DreamGarden, an AI-driven assistant for game design, which works semi-autonomously toward a high-level goal by growing a “garden” of plans and actions, where intermediary output is human-interpretable and editable. In designing this system, we pursue a sliding scale between autonomy and reactivity. Our target for a game design assistive tool is in the early ideation stages and prototyping, with a goal of eventually enabling rapid prototyping of playable game snippets. As a first step towards this, we explore a tool which elaborates on an initial game design plan and then proposes a concrete implementation plan for building the experience in the Unreal game engine, which it then delegates to various agents in an orchestration that is decided on recursively by an LLM planner. DreamGarden is a first such attempt at an automated game prototyping tool to our knowledge.
To validate the system and explore its potential as a new means of human-computer interaction, we conduct a study with the following research questions in mind:
Is the system capable of transforming open-ended, potentially dream-like natural language prompts into functioning 3D simulation environments?
Is the system’s hierarchical and iterative process intuitive and accessible to users?
Does DreamGarden’s user interface provide ample means for users to intervene in this process, at various points in time and at various levels of abstraction?
2. Related work
2.1. Open-ended systems and generative gardens
DreamGarden is conceived of both as a designer facing tool and a self-perpetuating ecosystem of design ideas and content. In terms of this latter autonomous nature, we take inspiration from the field of Artificial Life (Aguilar et al., 2014), which partially involves the design of simulated environments in which complex and life-like phenomena are able to emerge automatically. Dave Ackley’s Moveable Feast Machine (MFM), for example, offers a compelling early example of a decentralized computational model designed for infinite scalability (Ackley and Cannon, 2011). Drawing inspiration from its evolving, adaptable ecosystem, we conceptualize our system as a form of a “garden.” In this framework, we design agents focused on specialized tasks ranging from asset generation to code execution, including one that focuses on automated debugging akin to the self-healing wire within the MFM. Through the interaction of these agents, our system exhibits continuous growth fostering emergent behavior within game design.
More recently, the Artificial Life system Lenia (Chan, 2018), a continuous extension of John Conway’s ubiquitous Game of Life (Izhikevich et al., 2015) has been combined with open-ended search algorithms such as Quality Diversity evolutionary search (Pugh et al., 2016) for the automatic discovery of novel and interesting “life-forms” (Faldor and Cully, 2024).
Other approaches to open-ended search involve varying degrees of human involvement. (Charity and Togelius, 2022) develop an online Twitterbot which seeks to autonomously generate aesthetically pleasing gridworld game maps by soliciting human feedback via automated polls. Users are also able to design their own levels on the platform, which are pitted against output from a small generative model that learns to imitate highly-rated levels. PicBreeder (Secretan et al., 2011), on the other hand, is entirely dependent on user involvement. But given its idiosyncratic representation of images as Compositional Pattern Producing Networks (Stanley, 2007), the participation of users in interactively evolving generative designs results in the discovery of novel designs without users setting these discoveries as explicit objectives in advance. Though DreamGarden is framed as a system for translating a prompt to a corresponding end product, the plans and artifacts it sometimes generates (e.g. to generate 2,000 blades of grass individually, or the expressionistic and self-colliding meshes resulting from initial attempts to procedurally generate standard objects) could potentially be explored as art objects on their own merits, with the system’s intermediary outputs repurposed for divergent, open-ended search (Stanley and Lehman, 2015).
2.2. Generative design assistants
Large models have been increasingly studied for their potential in assisting in creative design tasks. The authors of (LING et al., 2024) developed a web-based tool which uses text-to-image models to help users prototype their design concepts, rapidly translating between text and images with an eye toward bridging the gap that can emerge between artists and designers when communicating across these modalities. LLMs have been used to power an evolutionary engine for collaborative game design in (Lanzi and Loiacono, 2023). Such LLM-driven tools build on prior work using more conventional methods. Sentient Sketchbook (Liapis et al., 2013), for example, uses novelty search to provide alternatives to human designs, and automates playability checks; while (O’Neill et al., 2009) develop an authoring tool for machinima, automating aspects like character control and shot framing to help novice creators readily apply conventional cinematic forms.
LLMR (Torre et al., 2023) is another example of a multi-agent system that translates input prompts to code. Their platform is focused on the Unity game engine whereas we target more high fidelity game design in Unreal. Fundamentally, LLMR is a tool for casual creators (Compton and Mateas, 2015) enabling real-time scene and environment generation in VR for immediate use. We focus on game creation over an extended time horizon necessitating the use of hierarchical planning. Our tool offers more specialized usability for game designers, enabling them to begin with minimal high-level input followed by an automated and adaptable process. Whereas in LLMR, interaction with the system is designed to be seamless, with intermediary prompts and code largely concealed from the user, our user interface is designed to expose this process in an intuitive way, allowing ample opportunities for designer intervention.
2.3. World models
In reinforcement learning, where game-playing agents are trained via a reward corresponding to in-game score or success, world models—generative models of game mechanics and levels (Ha and Schmidhuber, 2018)—have been developed to alleviate the bottleneck of running expensive game simulations during agent training, often leading to highly performant game-playing agents (Hafner et al., 2020). Generally, these models predict the next frame of the simulation given the previous frame(s) and current agent action. Similar models have also been explored for their purely generative capability, with (Bruce et al., 2024) training a model on a large internet corpus of videos of gameplay of diverse games to predict next frames and infer player actions, then prompting it at runtime to generate next frames of imagined games given out-of-distribution input such as human-drawn sketches. (Valevski et al., 2024) showed that diffusion models can serve as game engines, with a demonstration of stable simulation of the classic game Doom running at 20 fps on a single TPU, though for relatively short times before losing stability.
One possible issue with such approaches, when thinking of assistive tools for game designers, is that they are trained to generate the game at the level of pixels. A design workflow involving such models, then, limits designer intervention to the provision of the initial prompt: the model does not manipulate any symbolic representation of the game world, and designers cannot tweak its output with any level of detail once the generation process is complete, nor do the models make use of existing methods for rendering assets or encoding game logic which might make their output more reliable and their generative task more tractable (as a result, these models often lose coherence quite rapidly after generating beyond the initial prompt).
The system developed in this paper, by contrast, develops an interface between generative models and an existing game engine, namely Unreal Engine (Epic Games, [n. d.]), a robust, open-source editor for game levels, assets and logic, popular both with indie developers and triple-A studios.
2.4. Automatic game generation
Prior work has explored the use of language models to generate games from scratch. Word2World (Nasir et al., 2024) generates a narrative from scratch—which tend to be of a Brothers Grimm fairy tale variety (Grimm and Grimm, 2004)—and subsequently uses this narrative to specify a set of characters, a tileset, and a sequence of grid-based level layouts. Individual tiles are then generated using the text-to-image model Stable Diffusion (Rombach et al., 2022). Here, game mechanics and objectives are essentially fixed in advance, and limited to player movement, the impassibility of certain tile types, and achievement of certain objectives by reaching special tiles.
By contrast, DreamGarden is restricted to generating 0-player simulations, though the mechanics—in terms of interactions between abstract Actor classes written in C++ —are unrestricted and can be determined from scratch by our code generation submodule.
The procedural generation of entire games has been an area of research prior to the emergence of large pretrained models, with (Cook et al., 2016) raising pressing questions around the valuation of creativity and the participation of such autonomous agents in game development communities. While our system could also be let loose as an autonomous agent, this angle is at present somewhat less viable because of the scope of our ambition in generating fully open-ended 3D simulations via arbitrary code. We also take for granted that—no matter how much the frontier of automatic game generation is pushed forward by the orchestration of increasingly sophisticated specialized models—the components of such a system will inevitably be repurposed as designer-facing tools, whether by experimental AI artists or seasoned developers. In this paper we therefore focus mainly on the possible modes of interaction afforded within such systems.
Rosebud AI (AI, 2023) serves as a notable startup in this domain, providing an AI-powered platform for creating browser-based games in JavaScript, along with a collaborative community where game developers can share and build upon each other’s creations. Our method offers a similar natural language-to-code tool but focuses on structured, plan-to-action workflow. This workflow guides users through the necessary sub-tasks to fully realize their creative vision. Furthermore, while Rosebud AI focuses on 2D browser games, we specialize in the 3D domain allowing us to support more complex asset creation and immersive environments.
2.5. Large models for code, asset and environment generation
Alongside the a nascent push to leverage large models for automatic game generation, LLMs have also been applied to game design sub-tasks, such as the generation of levels in Sokoban (Todd et al., 2023) and Super Mario Bros. (Sudhakaran et al., 2024). Meanwhile, large text-to-image models have been leveraged for the generation of structures and environments in MineCraft (Earle et al., 2024).
A large body of work has been focused on creation of assets that could be used in games. Much of this focus has been on the automated creation of 3D assets and textures, for example (Richardson et al., 2023; Chen et al., 2023; Liu et al., 2023; Faruqi et al., 2023; Tochilkin et al., 2024). Text2AC (Sun et al., 2024) is a hybrid approach using LLMs and diffusion models to generate game-ready 2D characters. Automated animation of rigged assets has been considered for example in (Huang et al., 2023) using LLMs to generate structured outputs of joint rotations at key frames. Animation generation has also been explored in (He et al., 2024).
OMNI-EPIC (Faldor et al., 2024) uses language models to generate 3D environments and tasks for embodied agents. DreamGarden similarly leverages large models for feedback on generated environments, but also considers aesthetic and non-functional aspects of the environment. Instead of being geared toward RL agents, DreamGarden is at root a user-facing tool. In future work, however, it could be combined with open-ended player agent learning to result in more functionally complex games with sophisticated player mechanics.
For a general overview of language models such as GPT in games, including in procedural content generation, we point the reader to surveys on the subject (Yang et al., 2024; Gallotta et al., 2024). Perceptions of game developers on the usage of generative AI tools in game design have been recently studied in (Boucher et al., 2024; Panchanadikar et al., 2024).
3. Methods
3.1. Design goals
We have three main goals in the design of our system. First, we explore the capabilities of large models in generating interpretable, designer-editable representations of functional game simulations from a single, high-level prompt. Second, we develop a user interface for intervening with the system at various points in the generation process and at varying levels of abstraction. Finally, we seek to validate our approach, by conducting an in-person usability study in which participants with interest or experience in game design interact with and give feedback on the system.
In order to effectively generate entire simulations from a single prompt, we argue that three elements are crucial. First, just as most game design involves multiple specialists (or at least the donning of many caps by a generalist designer) multiple specialized AI agents are required to competently perform the plethora of semi-independent tasks involved in full-fledged game design. Second, these specialists must be orchestrated dynamically and automatically, as not all game design endeavors involve the execution of the same tasks in the same order. Finally, this orchestration must be carried out via hierarchical planning, with high level goals broken down recursively into sub-goals and implementation tasks, to ensure coherence between high-level objectives and actual implementation steps.
In terms of the user interface, we develop a node-based visual interface for editing steps in the hierarchical planning tree, pruning or expanding sub-plans, and providing feedback on the simulation resulting from generated code and assets.
Via our usability study, we seek to answer our questions about the system’s capabiltities (Q1), users’ perception of its planning and implementation process (Q2), and their means of interaction with it (Q3). By recruiting a mix of professional and novice designers, along with more casual creators, we are particularly interested in the diversity of modes of interactivity that might emerge among users of the system, and how future development of the system can serve to accomodate them effectively.
3.2. System overview
In this section we give a detailed description of the DreamGarden pipeline.
3.2.1. Planning module
We begin with the planning module, whose input is the seed prompt, which is open-ended and may describe a dream, an imagined scenario, or a rough game sketch. This is fed into the broad planner, which is system-prompted to take this open-ended input and reformulate it as an outline for a video game to be implemented in Unreal Engine. This submodule is also asked to create a broad, high-level plan for implementing the game. This high level game description and broad plan is parsed, and converted to the root and child nodes—the game reformulation and broad plan steps, respectively—of a hierarchical plan tree.
Next is the sub-planner. This submodule is shown the entire plan tree in its current state, and asked to re-articulate a particular node (a step or sub-step of the plan) in more detail, then further expand it into a list of sub-steps. The sub-planner is applied to the tree breadth-first, and at each step its output is parsed, and a node is added to the working tree for each new generated sub-step. Both planner sub-modules are given a description of the available implementation submodules—each specializing in a particular concrete code or asset generation task—and told that the tree they generate must ultimately terminate in leaf nodes that each correspond to an individual task for one of these submodules. The sub-planner is able to flag a generated node as a leaf node by appending special text to its output that indicates to which submodule the node is to be assigned. The sub-planner’s tree generation process terminates when all plan nodes have children or are marked as leaf nodes.
The leaf nodes are then translated into prompts or other formatted inputs for the implementation submodules by the task generator, whose prompt is adapted to give general guidance on how best to prompt for the particular implementation submodule at hand. The corresponding implementation tasks are then executed in sequence—visiting sub-plan steps in order—with the generated assets from prior tasks fed as input to subsequent submodules. This process of recursively growing the plan tree then translating its leaf nodes into implementation task prompts is shown in Figure 3.
3.2.2. Implementation submodules
We now describe the specialized submodules designed to carry out individual implementation tasks.
Code generator
The coding submodule is the most general and complex submodule in DreamGarden (see Figure 4). It involves the generation of C++ code for Unreal Engine “Actors”—essentially, general discrete entities within the simulation—and the generation of an initial layout for these actors in the level, which layout is then instantiated via a hard-coded python script inside the editor. In general, we observe that code generation models like (Zhao et al., 2024; Lozhkov et al., 2024; Bai et al., 2023) are trained on large volumes of C++ code which grant them capabilities in implementing a wide range of functions in Unreal Engine. However, these models still require precise prompting and a well-structured system architecture to generate fully functional environments within Unreal. In our experiments, we use GPT-4o (OpenAI, 2024b) for its speed and versatility.
The task prompt for the coding submodule involves distinct actor and spawner prompts. First, the actor prompt is fed to a code-generator LLM, whose system prompt contains some directives on writing C++ actors for Unreal Engine, and lists of all materials and meshes available in the Starter Content directory, as well as a list of any meshes that have already been generated during asset downloading or generation tasks that have already been carried out. A set of complete C++ files are parsed out from this model’s output.
These files are then copied to the UE project directory, and the project is re-compiled. The compilation log (along with the generated code and task implementation prompt) is fed to a compilation evaluator submodule, which outputs a verdict of success or failure, providing feedback on the source of the compilation errors and guidance on how to resolve them in the latter case. If compilation is successful, then a spawn layout generator LLM is queried for a layout which places instances of generated actors in the scene by specifying their coordinates, scale, rotation, and initial values for any editable properties of these actor classes in the generated code. This LLM is shown the generated code, its portion of the implementation task prompt, and some relevant pointers such as the correspondence of UE units to centimeters. A layout json is parsed out from the output of this model, and the UE Editor is launched. A hard-coded python initialization script is run when the level is loaded, which places instances of the generated actor classes in the scene. Should this python script error out—in particular if the layout generator has referenced a class that does not exist in the project’s generated C++ code—then the relevant section of the UE log, the initial code, and task prompt are fed to an LLM to produce feedback and pointers on how to address the issue. Supposing the python script does not raise any exceptions, then actor instances are initialized in the scene and simulation code is launched. If the editor crashes due to runtime errors, the crash log is again fed to an LLM alongside code, layout and task prompt to produce feedback and guidance. If the simulation runs successfully, then a hard-coded ‘FilmCamera’ class (protected from being overwritten by generated classes) captures 6 screenshots of the first 6 seconds of in-engine simulation, and these screenshots along with the runtime log, actor code, layout, and task prompt are fed to a Vision Language Model (VLM) which must then output a success/fail verdict, and provide feedback in the latter case.
Supposing failure at any of the above steps, the resulting feedback, along with the generated (broken) code and layout, and the task prompt, are fed back into the code generator, repeating this process for a fixed number of attempts before the system is forcibly moved on to the next implementation task.
Procedural mesh generator.
A variant of the coding submodule, the procedural mesh generation submodule, involves modifying the coding submodule’s system prompt to request code for the generation of a procedural mesh for a given structure or terrain. An example of a procedural cube mesh written inside a C++ UE actor is included in the prompt as an example. If the code for the procedural mesh compiles successfully, a default layout is generated, wherein a single question of the actor in question is instantiated at the origin for visual examination. The same feedback loops, with a fixed number of attempts, are applied in this submodule. In Figure 2 we show outputs resulting from the chaining together of two procedural mesh generation tasks, with the first generating hilly terrain, and the second adding vegetation, which it adjusts to make more realistic and diverse after visual feedback.
Diffusion mesh generator.
The diffusion mesh generation submodule (Figure 5) generates novel 3D assets from scratch, by chaining together a text-to-2D diffusion-based image generator and a 2D-to-3D textured mesh generator. The task prompt for this submodule is a detailed description of the desired asset. This description is fed into a text-to-2D image generator. The resulting image is then fed into a 2D-to-3D generator (we use TripoSR (Tochilkin et al., 2024)).
The resulting 3D object file is then converted into an UE asset in the engine, and added to a folder within the game’s project, to be later exposed to coding submodules for use in the scene via generated actors.
The initial prompt is appended with some further directives in an attempt to make the image a suitable prompt for a 2D-3D model (i.e. requiring the object to be fully visible and posed against a blank background).
Mesh downloader.
The mesh download submodule (Figure 6) uses text descriptions to retrieve and import handmade assets from a pre-existing database. For this submodule, the task generator produces a concise description of the desired asset. This text is then embedded via CLIP (Radford et al., 2021). The CLIP embeddings of uploaded thumbnails of all assets in the Objaverse database (Deitke et al., 2024) are precomputed, and the distance between these thumbnails and the generated text description are computed, and the the asset whose thumbnail has least CLIP distance to the text description is downloaded from the database.
Both the asset downloader and diffusion mesh generation submodules launch UE in the final step, and, via a hardcoded python script, import the resultant GLB or gLTF asset files into UE, converting them to textured meshes within the engine and saving them to disk for later use.
3.2.3. Graphical user interface
We develop a Graphical User Interface (GUI) for DreamGarden, to expose its growing the user’s seed prompt in a more intuitive and visually informative way. Because the action plan is effectively a tree, the basis of our GUI is a node-based editor, i.e. an interface in which users can manipulate nodes in a tree. Nodes are either (sub-)steps in the hierarchical plan, implementation tasks, or steps taken by implementation submodules, and edges correspond to the hierarchical relationship between planning steps and tasks, and the the chronological flow of implementation steps. The GUI is shown in Figure 7
To facilitate responsiveness between the system’s backend and the GUI, we structure the backend call as a repeated function on a “frontier” of nodes to be expanded/implemented. The user’s edits can effectively result in pruning this tree, modifying nodes, and adding/removing them from this frontier.
The current user workflow is as follows. The user begins with an empty canvas, and can add a seed node, where they enter the open-ended high level prompt. They then have the option to “step” or “play” the system, either stepping through a single call to the backend node-expansion function, or initiating the repeated expansion of nodes. (The inclusion of the step function is crucial in cases where users intend to take a more active role: if instead the backend is set to play, then user edits on nodes that have already been expanded may result in the computationally costly expansion of nodes which are ultimately be pruned as a result of user intervention.)
The seed node is expanded first into a broad plan, with children nodes containing plan steps fanning out to the right. This kind of expansion is applied recursively to the tree, until all extant nodes have children or are leaves. The placement of nodes can be adjusted, and they can be expanded to show the full text generated by the system which describes the plan step they represent.
Once the plan tree is fully expanded, the leaf nodes of the tree are added to the frontier and queued for transformation into task prompts. That is, the task generator will iterate through these leaves in order, transforming them into a specific prompt (or other input) for the submodule to which each task is assigned, which prompt is represented in a new node to the right of its corresponding leaf plan node. While all plan nodes are represented in neutral light grey, these task nodes are
User edits
Each node in the action tree has an “is leaf” checkbox. When toggled by the user, this determines whether or not the given node is considered as a leaf in the tree (i.e. the description of a concrete implementation task to be handed off to a particular submodule). By turning a non-leaf into a leaf node (i.e. when the system has unnecessarily expanded a relatively atomic step into sub-steps), the user effectively prunes the tree. During the next call to the node-expansion function, any plan nodes that are (grand-)children of this node are removed. If any extant children are leaves that are attached to task nodes, these task nodes are additionally removed from the tree. Similarly, if these tasks have already resulted in implementation step nodes, then these implementation nodes are also removed (this data is backed up to a separate folder, should the user want to reverse their changes). Finally, because any code and assets resulting from these implementation steps would have been handed down to implementation tasks following it in the ordered sequence of leaf nodes, then any implementation steps belonging to these tasks are also backed up and deleted.
By turning a leaf node into a non-leaf node, the user effectively adds this node to the frontier, queuing it for expansion into a sub-plan. Again, if a task node was generated for this leaf, it is deleted, along with any implementation nodes belonging to this task, or any implementation nodes belonging to tasks following it in the ordered list of tasks.
During code (and procedural mesh code) generation, coding attempts and code evaluations are displayed in nodes that are chained in a sequence, moving from left to right out from the parent task node (itself stepping from a leaf node of the plan tree). When the system produces code that compiles and runs successfully in Unreal, screenshots are automatically captured of the first moments of simulation. These screenshots, along with feedback produced by a vision language model based on them (in addition to the runtime logs and generated code and actor layout), are displayed in a special “visual evaluation node”, which is chained to the “code attempt” node from which it stemmed. These nodes are augmented with a “compile & run” button, which allows the user to compile the code as it was at this state, and launch the editor with the corresponding actor layout. The user is then able to inspect the resultant simulation in Unreal Engine. They may, for example, make edits to generated code (testing them via recompilation from within the editor), which when saved would then be fed into subsequent iterations of code generation. They may also return to the GUI, and edit the feedback generated by automatic evaluation. This will then result in the pruning of all implementation steps having this node as parent (or, spatially speaking in the GUI, to the right of or below it), and the rerunning of the subsequent code generation iteration following it. In Figure 8, for example, the user has requested a self-playing city simulation. The user launches the code at a state wherein the code generator has procedurally laid out generated meshes in a grid-like pattern. Via in-editor observation, they confirm the absence of any kind of ground plane beneath the city layout. Entering this feedback into the relevant GUI node and resuming the system results in a subsequent node where terrain has been added to the scene.
3.3. Study design
To explore the interaction affordances of our system, validate it against user needs and behaviors, and explore directions for future development, we conducted a usability study with participants. These participants were were recruited via emails and posts in messaging groups from within the researchers’ organization. Users first filled out an intake survey, in which they specified their levels of experience with game design, generative AI, and Unreal Engine. The users were then given a brief overview of the system’s intended purpose, and the means they had available to interact with it. Users were first invited to imagine a high-level seed prompt, with little restriction on form or content, which they then entered into the initial “dream” node in the GUI. The study was conducted in person, with the DreamGarden system running on a laptop with a top-tier GPU. Users were encouraged to observe the system’s operation by interacting with the GUI, as it expanded their prompt into an increasingly detailed plan of action, assessing its viability by talking through it out loud. They had the option of pruning or expanding the tree further by toggling leaf node status via the GUI. They were then invited to observe the operation of the implementation submodules; examining assets in the GUI or a separate file explorer window, and assessing generated code files in a separate IDE window if desired, and to the best of their ability. Users were also reminded at this point that, as per the consent form they had signed prior, they were free to check their phone, grab a drink, or otherwise disengage from the system’s operation as it grew artifacts autonomously. When the system generated code that compiled and ran successfully in-engine, they were invited to observe visual feedback nodes, compile and run the code at points that seemed interesting, explore the resulting environment via the UE Editor and provide customized feedback in the corresponding GUI nodes based on their first-hand observations. Finally, after using DreamGarden, the participants were asked to answer a series of questions in the exit survey on their experiences with it, including rating their frustration and enjoyment of using the system on on a 5-point Likert scale (1 = Strongly disagree, 5 = Strongly agree). See the Supplementary Material for the detailed questions. Studies lasted a maximum of 45 minutes each. The participants were invited to submit seed prompts ahead of arrival, allowing them to step ahead at their own pace through the generation of a partially complete garden without having to wait for the system to catch up.
The user study design, including calls for participation, the structure of the in-person study, the user workflow with respect to the system, and the intake and exit survey questions, was approved by an Internal and Ethics Review Board prior to conducting the study.
3.3.1. Data collection and analysis
In addition to the intake and exit surveys, the researcher present during the study took notes of the participants experience with and commentary on the system. Additionally, each participant consented to audio and screen recordings of their study session. Audio recordings captured participants’ comments during their interaction with the system and discussion surrounding the design process. These recordings were transcribed using Microsoft Teams transcription software. Screen recordings captured interactions with the system (both the GUI as well as the file explorer and code editors used by some users to more closely track the system’s output). Finally, DreamGarden automatically logged the work of the system in concert with the user—including any user edits, and backups in the case of any resultant tree-pruning—that is, the sequence of all prompts, responses, generated code and artifacts, and in-engine screenshots produced over the course of the study.
4. Results
4.1. Qualitative system observations
We experiment with various hyperparameters, most notably the branching factor and maximum depth of the tree generated by the planning module, and the maximum number of attempts allowed to the coding generator before forcibly moving on to the next task. Note that the branching factor and depth limits are insisted upon in the system prompts of the planning module, though they could also be manually enforced by throwing away supplementary children and forcing certain nodes to be leaves (and querying the language model separately to assign it an appropriate implementation submodule), respectively.
We note that the system will not always expand the tree to the fullest extent. Still, in simpler scenarios, like “a sheep grazing on a grassy hillside”, it will often superfluously subdivide certain steps, e.g. breaking down “generate sheep asset” into separate tasks for the generation of individual body parts, or for the generation of mesh and texture. Note that this latter case speaks to a problem we notice occasionally with the system, in which it will invent tasks beyond the capability of the submodules to which it currently has access, or at least which are not explicitly mentioned to it. This was the case with lighting, atmosphere effects (such as fog) and skyboxes, for example, where the system would sometimes implement these features despite their conflicting with the default level’s default lighting. In this case, we subsequently removed all environment assets from the default level, prompted the coding module to account for these explicitly, and gave an example of a sufficient ‘EnvironmentManager‘ Actor class in its prompt. With animation, on the other hand, we adapted our prompts to insist that the system would not be capable of animating skeletal meshes.
When we attempt to restrict the tree to particularly small depths and branching factors, it sometimes ignores these requests, for example adding too many children to the root node.
We note anecdotally that for more complex tasks, larger depths and branching factors allow for better subdivision of implementation tasks, making it easier for the system to progress toward high-level objectives without getting stuck on implementation tasks involving extending the code in very complex ways in one pass. Instead, it is easier for the system to add small but observable functional changes to the code step by step.
Naturally, allowing the code generation submodule more generation attempts increases the odds of it finding code that compiles and runs, and results in acceptable visual results. This improvement seems largely attributable to the specific feedback given by evaluation agents, which are often able to make sense of informative compilation and runtime errors and warning messages, suggesting relevant code fixes for future generation attempts.
4.2. Usability study results
We will now discuss findings from the user study and group them into a few broad categories. Our survey indicated that all the users were at least partially familiar with generative AI, though their familiarity with game design tools was more varied, while Unreal engine itself was largely unfamiliar to most of the participants, see Figure 9. We present sample intermediary results from the user study generations in LABEL:fig:figures.
Utility of planning
A common theme expressed among users was the utility or potential utility of DreamGarden in the planning and prototyping phases of design, or for creating first drafts of more complex projects. Indeed, several users expressed that the planning module may have some standalone utility, irrespective of the system’s ability to execute these plans itself. Conversely, users expressed reservations regarding the potential use of the tool in cases demanding more complex game mechanics and environments. Though not prompted to directly consider the merits of DreamGarden as an educational tool, a natural split on this matter emerged among some users with P10 stating that “seeing a game design being broken down into steps may be useful to new developers trying to learn how to build their own game”, and P3, on the other hand, stating that they “would be hesitant about using it for the purpose of really learning, since it takes some of the opportunity to problem solve away from the user”. Both P3 and P10 were familiar with game design and generative AI, and very unfamiliar with Unreal Engine.
When not to use such a tool
When asked to consider what circumstances were not befitting of the application of a tool like DreamGarden, several users imagined scenarios in which designers desired higher degrees of control, or more fine-grained control over the final product, with P4 stating “I would be reluctant to use a tool like LucidDev111Note that the working title of the system at the time of the creation of the user study survey was “LucidDev”, and was later changed to DreamGarden. for any kind of deliberately artistic endeavor, as I would prefer to have direct control over the outputs”. P9 raised the issue of the potential limitations of language, stating that the system would be unsuitable in situations “where I don’t have words to describe what I want”; as well as scenarios in which measures of success were not necessarily visually obvious, or easily described in advance, but rather required “some intuition to decide if it is good enough”.
Enjoyability
When asked to discuss elements of the system and experience which they enjoyed or did not enjoy, several users highlighted the pleasure of watching the system grow their seed prompt into a tree-structured plan action, with P8 stating that “the tree-like visualization of progress and what is the system doing at every step is pretty remarkable” and P2 stating that they “found it exciting to track how the system interpreted my simple prompt and how I could see it’s plan”. The fact that this relatively fast pace is not maintained in the GUI (with, e.g., intermediary steps of the coding submodule being hidden from the front-end) seems to have caused some displeasure, with P2 stating that “the slow process of generating the scene wasn’t very enjoyable”. In this same vein, some users expressed interest in having more intermediate states visualized in the GUI, with P8 suggesting that “the code output can be shown on the browser window too so you have an even more detailed description of what’s going on”. Some of the intermediary output of the system generated over the course of the user study is shown in Figure 10.
Users expressed a range of emotions with respect to their experiences with the system Figure 11. In their explanations of these ratings, it is clear that expectations and investment play a large role. While P4 experienced high frustration (4) because the system “doesn’t really work”, they also experienced high enjoyment (4) because “it was fun to try out and see what it was doing”. P3, meanwhile experienced low enjoyment (2) and frustration (2), explaining that they were “not super invested in the outcome”.
Transparency and Explainability
In one notable example, P8 requested a “Super Mario Bros.-like” platformer “but with the art style of MineCraft” in their seed prompt. However, the system lost the intended MineCraft influence as it sub-divided the plan, generating generic NPC assets instead of stylized ones. Thanks to the explicit nature of all steps in hierarchical plan generation via the GUI, the user was able to pinpoint the expansion step at which this request was ignored. This would have made it straightforward to remedy via human intervention. Cumulative errors in code, on the other hand, were not obvious to users, who generally wound up disoriented and lagging behind system progress when digging through generated code and evaluation—here, output is verbose, dense, and less accessible through the GUI. Along these lines, several users suggested providing summaries of code changes during generation, or errors during evaluation, with reference to relevant code snippets, making the system’s functioning “easier to digest” to users, such that they might “jump in and give it some correction so it doesn’t have to flounder so much on its own” (P7).
Interactivity
While some users seemed to envision a more generally capable and fully autonomous system, capable of making functional games from scratch without intervention, others imagined ways in which the system might depend on human designers, and even explicitly ask them for help given particularly difficult subtasks or uncertain results. P8, for example, thinks it “less important that I give it a simple prompt and come back the next day hoping it did the right thing, and more important that I can craft and sculpt with it directly”, for example “I could be like ”make a big cube in the center” and *poof* a big cube appears”. P2, meanwhile, thinks “more iteration on each step would make the end result better”, and that “this could be human in the loop or not”.
P4, who was slightly familiar with Unreal, spent time investigating generated source code, using file editors to examine the sequence of prompts and responses produced by the coding submodule.
Though users were all reminded that in their consent forms, they were welcome to take breaks (check phones, grab drinks) during the study, due to the semi-autonomous and long-timescale nature of the system, only P3 actually left the room to get a drink from a nearby kitchen area. Several users, on the other hand, initiated some form of small talk with the researcher present while the system iterated on code.
Opinions on the GUI
P3 pointed out the fact that so-called “leaf nodes” in the plan tree are seemingly not actual leaves in the overall graph as displayed in the GUI, given that they themselves appear to be the parents of task nodes. Similarly, P5 expressed repeated confusion about the togglable “is leaf” checkbox available in plan nodes.
Most users focused on the GUI, with several asking for an explanation about the coloring of nodes. Some users (P2, P8) highlighted appreciation for being able to preview diffusion-generated assets in the GUI. P3 suggested adding highlighting to the in-progress node, and P3 and P7 both pointed out that coloring nodes red gave a false impression that this node had failed (in the current version of the system, nodes containing screenshots from successfully compiled code that has run in the engine, along with the feedback given based on these screenshots, are colored red, regardless of the overall verdict handed down in this feedback). P3 suggested reserving “special” colors like green, yellow and red, for respective states of success or failure.
When the system writes functional code, the editor is ultimately launched to capture screenshots of the simulation at this stage—but only 6 screenshots for six seconds. Often, a user would exclaim when, sometimes after a long lull of the system silently iterating on faulty code, an editor window with something in the distance appeared out of the blue. But of course this window was then automatically killed almost instantly. This seemed to appear as a somewhat frustrating phenomenon for users, who were sometimes able to begin moving toward the centered object in the simulation in the very last available second.
In terms of hypothetical improvements to be made to the system, some shared desiderata include being able to critique and/or refine assets in the GUI, the ability to edit or delete arbitrary nodes in the garden, and swap out submodule assignment.
5. Discussion & Design Implications
In this section we discuss the findings from the usability study and the implications for design of future versions of assistive systems for game development.
Improving overall performance
Most obviously, the fact that the system struggled with more complex scenarios, as many users pointed out (or suspected, given the limited time allotted to the interactive portion of the study itself—around 25 minutes—relative to the time actually required by the system to finish iterating on a seed prompt given the default hyperparameters—around 1 hour) points to a general need to make the system more performant overall. We concede that large models may not be able to produce genuinely novel artifacts relative to the corpus of human-generated content upon which they have been trained. Nonetheless, by integrating them in a hierarchical, iterative, and specialized pipeline such as DreamGarden, we argue that they can at least provide a powerful means of interpolating among this vast space of prior human content, a source of what Kate Compton has termed “liquid art” (Karth and Compton, 2023). Perhaps the simplest way of making the system a better interpolator of existing artifacts would be to use more thorough few-shot prompting. Though there is a vast amount of documentation, tutorials and example code snippets pertaining to Unreal Engine online, we have largely refrained from incorporating it into our prompts in the current iteration of the system. One concern might be where to begin with such a sheer quantity of potential prompting material. But thanks to the sub-divided and iterative nature of design tasks endeavored by DreamGarden, some obvious opportunities for inclusion emerge. In particular, during the task generator’s creation of submodule prompts/inputs, it could additionally select what libraries’ documentation should be included in the prompt for the coding submodule with each task, given the description of this task in the leaf plan node, and the names (and perhaps brief descriptions) of high-level C++ modules in UE. Then, the API pages for these libraries could be inserted into the code generator’s prompt. Similarly, when the coding module encounters compiler or runtime errors, the corresponding evaluator LLM might choose to include relevant API documentation in the prompt for the code generator at the following attempt, to rectify any incorrectly called functions or classes.
More digestible presentation
Regarding the possibility of providing more digestible code snippets, we note that, though in some cases this might involve a tradeoff between interpretability and computational expense—wherein models would have to prompted, after generating or evaluating code, to provide summaries of their work—our generators and evaluators are largely already being chain-of-thought prompted (Wei et al., 2022) to some degree. Code evaluators, in particular, usually summarize issues with the code in numbered lists of brief bullet points, which are elsewhere in its response expanded in more detail. This kind of structured response could be explicitly requested and parsed out to be emphasized in relevant GUI nodes, with the full text expandable via user interaction. Code generators, meanwhile, sometimes provide brief summaries of their work in natural language before or after the prompt. We expect that asking for more structured outputs would come at little extra cost to the quality or efficiency of their work and might even make it more robust.
Broader applicability
The general enjoyment of users had in watching their plans expand suggests that this approach developing high-level into low-level prompts may have broader applicability. In particular, it may be that for many applications, breaking a high-level goal down into a hierarchical plan might serve to improve the robustness of generated content (at the price of additional compute). We emphasize that the tree-based planning approach is agnostic to the exact nature of the submodules to be called upon by leaf nodes. One could thus apply the system to, e.g., robot control problems, the creation of abstract multimedia art, or the generation of a complex narrative.
One might ask how exactly a hierarchical planning tree involving iterative breadth-first expansion via repeated model queries differs from a hierarchical planning tree generated in one pass. The key difference is that we can control exactly what is provided to the context window during expansion of each node based on said node’s placement in the tree, whereas, during one-pass tree generation, any approach will, for example, ultimately expose the last node to all prior nodes, because of the linear nature of the text being generated222We liken this, by analogy, to the “graph convolutions” in neural network architecture (Wu et al., 2020), which structure the signals passed to neurons based on a graph structure..
Here, more generally, we see hints of a possible alternative to the standard, linear “chat” interface, where we instead represent the open-ended generation of text as a graph, where this graph serves to visualize what (AI- or user-)generated text ought to influence each new generation, with users recombining fragments of text in a way that is more free-form and intuitive than that afforded by strictly linear walls of text. Users’ pleasure in interacting with DreamGarden’s node-based GUI would seem to bode well for such a vision.
Especially for open-ended, creative tasks, it would seem that a graph-based approach, resembling the classic “mind-map”, involving an ever-enlarging frontier of ideas moving out across an infinite plane, serves as a better symbolic representation of AI-augmented, human-driven discovery. Though AI can make “liquid” via sophisticated interpolation the space of prior human expression, by incorporating humans in the open-ended loop one can begin to probe the frontiers of this space, expanding and reshaping it, stretching it to fit new, novel artifacts.
Question of control
Regarding the concern about degrees of control in situations where the authors views themselves as “auteur”, we note that in theory, the system could play a much more passive role than it does currently. For example, the system could be paused by default. A designer could then produce any number of C++ actor files, assets, and create a scene layout, only relinquishing control to the system when, for example, they wanted to receive feedback on how to address a compilation or runtime issue. Similarly, the user could provide a partially complete project, and experiment with “modding it”, by giving over control for subsequent tasks intended to extend the current project. Or, conversely—and as was a common suggestion among users—the system might be used to generate a rough first draft or prototype which is then taken over entirely by a human designer.
Though this promise—of a system with, effectively, a knob that adjusts the degree to which it is autonomous vs. reactive—is seemingly grand, and risks stretching itself thin from a design standpoint by aiming to please too many users too much of the time, it should be relatively easy to deliver if we can ensure that the system is able to observe all edits made by the user, and, conversely, the user is able to observe all edits made by the system. It need not necessarily be the case the the assistant and the user have exactly the same means of expression (and this is likely too lofty of a goal), but rather that any action on the part of either can be entered into the conversation. This is sufficient to allow the system to transition from complete autonomy to complete reactivity, because it ensures that no matter whether user or system take over completely for any amount of time, the other is always able to observe and respond to the changes that have occurred over this period. To this end, the main focus moving forward with the development of the system should be twofold: both to increase the legibility of the system to the user, for example by following some of the practical suggestions pointed to by users in our study; and to expose more elements of the UE Editor to the system.
Direct manipulation in the editor
Though the system can currently observe edits to code made by users, it could also conceivably respond to modifications made to the placement and other features of objects within the editor. Currently, the system generates the initial layout (in terms of position coordinates, scale, and rotation) and initial parameters (e.g. the initial speed or direction of a moving object) of actor instances by producing a structured json containing this information. Functionality could be added to record changes made to these parameters, or the addition/deletion of actors, by the user to the level during their interaction with the editor, such that the next iteration of code generation would take this updated layout information as input.
We note a general frustration with the inability to directly observe generated artifacts in the UE editor. This is largely an engineering challenge that could conceivably be overcome in future iterations using one of several approaches. Most notably, we are currently not compiling in UE’s “LiveCode” mode from our python backend, though this is possible in theory (instead, we are forced to compile with the UE Editor closed, then relaunch it each time new code is generated). This change would allow the editor to remain open, even as our system was in the process or editing C++ Actor classes or adding new textured meshes to the scene. This could allow for a more fluid “dialogue” between human users and the system, in future iterations, with the user potentially making changes to generated code or layouts inside the Editor, then having the system iterate on the project while taking these user edits into account.
In addition or separately from seeing newly generated artifacts reflected in the Editor, the user may also want the option of exploring partially completed environments while at the same time letting the system “work ahead” (on later iterations of the environment or on later implementation tasks). Currently, the option to “compile and run” partially-complete environments, presented as a button on visual feedback nodes in the GUI, requires pausing the backend, since otherwise, collecting visual feedback at future steps, which terminates by manually killing UE processes, would result in killing the instance of the engine intended for the user. This could likely be addressed by tracking which process ID belongs to different instances of the Editor, and killing only the appropriate processes from the backend (this is complicated somewhat in cases where the Editor has crashed, but it can perhaps be assumed that the user has not crashed their instance of the engine at the same moment as the backend’s instance).
GUI improvements
Uncertainty about node colors demonstrates the need for a legend, or for more explicit node type labelling. Further, colors like green, yellow and red should likely be reserved to indicate (partial) success/failure of artifacts generated by the coding submodule, which is able to iterate on errors in its output. The system’s operation could also easily be made more transparent by highlighting the node currently in-progress.
6. Conclusion & Future Work
DreamGarden is a prototype of a tool that shows potential in being a useful assistive tool for game developers. The idea of growing a plan of actions from the initial prompt, however, is much more broadly applicable to game design beyond development in Unreal engine. Firstly, game developers use many tools beyond the game engine, and those could be exposed to the system as intermediate stages before attempting to build the experience in Unreal. For example, part of the initial implementations and prototyping could be done in a 2D setting, while some visual effects and procedural content could be developed in Houdini (SideFX, [n. d.]).
More generally, the ideas presented here can be applied outside of game development and apply more broadly in software development. In particular, the idea of automatically growing a tree of plans that can be implemented by particular agents seems like a simple solution to deciding an orchestration of multi-agent systems tailored to a particular problem, assuming that the recursive application of the planner can discover a reasonable plan of action. This suggests at least the necessity to fine-tune an LLM on proven high level implementation plans of existing projects, both in the gaming industry and more broadly in the realm of software development.
Regarding our initial research questions, we find that, while the system often struggles to autonomously generate complex simulations (Q1), users find substantial value in the intermediary planning steps and artifacts it generates in the process (Q2). Despite user studies being limited to a relatively small amount of time, we note the emergence of diverse modes of user interaction with the system (Q3). Inspired by the experiences and feedback from users in our usability study, we also conceptualized various means of fleshing out these modes of interaction to bolster DreamGarden’s co-creative potential.
In particular, we note the strong appeal of such a branching approach to “growing” content collaboratively with AI agents, and argue that a graph-based approach is a natural fit for creative and open-ended tasks. We also see that, even given our relatively rudimentary GUI, the paradigm of “gardening” as opposed to constant active human guidance or a fully passive “prompt and sit back” approach, invites a variety of modes of interaction from users with differing intents and levels of experience. The key here is in combining autonomous feedback and self-improvement with a variety of means of user intervention. This is strengthened by our focus on having the system operate on human-interpretable representations of generated content inside an established game engine, in contrast to extant efforts that seek to replace game engines with, effectively, uninterpretable latent variables.
Acknowledgements.
We would like to thank Cyan DeVeaux and Amina Kobenova for their invaluable feedback and support during the development of DreamGarden and the writing process. We’d also like to thank Jaron Lanier and Judith Amores for their guidance and encouragement during the project’s development.References
- (1)
- Ackley and Cannon (2011) David H. Ackley and Daniel C. Cannon. 2011. Pursue Robust Indefinite Scalability. In 13th Workshop on Hot Topics in Operating Systems (HotOS XIII). USENIX Association, Napa, CA. https://www.usenix.org/conference/hotosxiii/pursue-robust-indefinite-scalability
- Aguilar et al. (2014) Wendy Aguilar, Guillermo Santamaría-Bonfil, Tom Froese, and Carlos Gershenson. 2014. The past, present, and future of artificial life. Frontiers in Robotics and AI 1 (2014), 8.
- AI (2023) Rosebud AI. 2023. First Look at Rosebud AI Game Maker: Type, Create, Play. https://www.rosebud.ai/blog/first-look-at-rosebud-ai-game-maker Accessed: 2024-09-11.
- Bai et al. (2023) Jinze Bai, Shuai Bai, Yunfei Chu, Zeyu Cui, Kai Dang, Xiaodong Deng, Yang Fan, Wenbin Ge, Yu Han, Fei Huang, Binyuan Hui, Luo Ji, Mei Li, Junyang Lin, Runji Lin, Dayiheng Liu, Gao Liu, Chengqiang Lu, Keming Lu, Jianxin Ma, Rui Men, Xingzhang Ren, Xuancheng Ren, Chuanqi Tan, Sinan Tan, Jianhong Tu, Peng Wang, Shijie Wang, Wei Wang, Shengguang Wu, Benfeng Xu, Jin Xu, An Yang, Hao Yang, Jian Yang, Shusheng Yang, Yang Yao, Bowen Yu, Hongyi Yuan, Zheng Yuan, Jianwei Zhang, Xingxuan Zhang, Yichang Zhang, Zhenru Zhang, Chang Zhou, Jingren Zhou, Xiaohuan Zhou, and Tianhang Zhu. 2023. Qwen Technical Report. arXiv preprint arXiv:2309.16609 (2023).
- Boucher et al. (2024) Josiah D Boucher, Gillian Smith, and Yunus Doğan Telliel. 2024. Is Resistance Futile?: Early Career Game Developers, Generative AI, and Ethical Skepticism. In Proceedings of the CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 173, 13 pages. https://doi.org/10.1145/3613904.3641889
- Bruce et al. (2024) Jake Bruce, Michael D Dennis, Ashley Edwards, Jack Parker-Holder, Yuge Shi, Edward Hughes, Matthew Lai, Aditi Mavalankar, Richie Steigerwald, Chris Apps, et al. 2024. Genie: Generative interactive environments. In Forty-first International Conference on Machine Learning.
- Chan (2018) Bert Wang-Chak Chan. 2018. Lenia-biology of artificial life. arXiv preprint arXiv:1812.05433 (2018).
- Charity and Togelius (2022) Megan Charity and Julian Togelius. 2022. Aesthetic bot: interactively evolving game maps on twitter. In Proceedings of the AAAI Conference on Artificial Intelligence and Interactive Digital Entertainment, Vol. 18. 18–25.
- Chen et al. (2023) Dave Zhenyu Chen, Yawar Siddiqui, Hsin-Ying Lee, Sergey Tulyakov, and Matthias Nießner. 2023. Text2Tex: Text-driven Texture Synthesis via Diffusion Models. arXiv preprint arXiv:2303.11396 (2023).
- Compton and Mateas (2015) Kate Compton and Michael Mateas. 2015. Casual Creators.. In ICCC. 228–235.
- Cook et al. (2016) Michael Cook, Simon Colton, and Jeremy Gow. 2016. The angelina videogame design system—part i. IEEE Transactions on Computational Intelligence and AI in Games 9, 2 (2016), 192–203.
- Deitke et al. (2024) Matt Deitke, Ruoshi Liu, Matthew Wallingford, Huong Ngo, Oscar Michel, Aditya Kusupati, Alan Fan, Christian Laforte, Vikram Voleti, Samir Yitzhak Gadre, et al. 2024. Objaverse-xl: A universe of 10m+ 3d objects. Advances in Neural Information Processing Systems 36 (2024).
- Earle et al. (2024) Sam Earle, Filippos Kokkinos, Yuhe Nie, Julian Togelius, and Roberta Raileanu. 2024. Dreamcraft: Text-guided generation of functional 3D environments in Minecraft. In Proceedings of the 19th International Conference on the Foundations of Digital Games. 1–15.
- Epic Games ([n. d.]) Epic Games. [n. d.]. Unreal Engine. https://www.unrealengine.com
- Faldor and Cully (2024) Maxence Faldor and Antoine Cully. 2024. Toward Artificial Open-Ended Evolution within Lenia using Quality-Diversity. arXiv preprint arXiv:2406.04235 (2024).
- Faldor et al. (2024) Maxence Faldor, Jenny Zhang, Antoine Cully, and Jeff Clune. 2024. OMNI-EPIC: Open-endedness via Models of human Notions of Interestingness with Environments Programmed in Code. arXiv preprint arXiv:2405.15568 (2024).
- Faruqi et al. (2023) Faraz Faruqi, Ahmed Katary, Tarik Hasic, Amira Abdel-Rahman, Nayeemur Rahman, Leandra Tejedor, Mackenzie Leake, Megan Hofmann, and Stefanie Mueller. 2023. Style2Fab: Functionality-Aware Segmentation for Fabricating Personalized 3D Models with Generative AI. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology. 1–13.
- Fielding (2012) Nigel G Fielding. 2012. Triangulation and mixed methods designs: Data integration with new research technologies. Journal of mixed methods research 6, 2 (2012), 124–136.
- Gallotta et al. (2024) Roberto Gallotta, Graham Todd, Marvin Zammit, Sam Earle, Antonios Liapis, Julian Togelius, and Georgios N Yannakakis. 2024. Large language models and games: A survey and roadmap. arXiv preprint arXiv:2402.18659 (2024).
- Grimm and Grimm (2004) Jacob Grimm and Wilhelm Grimm. 2004. The annotated brothers Grimm. WW Norton & Company.
- Ha and Schmidhuber (2018) David Ha and Jürgen Schmidhuber. 2018. World models. arXiv preprint arXiv:1803.10122 (2018).
- Hafner et al. (2020) Danijar Hafner, Timothy Lillicrap, Mohammad Norouzi, and Jimmy Ba. 2020. Mastering atari with discrete world models. arXiv preprint arXiv:2010.02193 (2020).
- He et al. (2024) Kevin He, Annette Lapham, and Zenan Li. 2024. Enhancing Narratives with SayMotion’s text-to-3D animation and LLMs. In ACM SIGGRAPH 2024 Real-Time Live! (Denver, CO, USA) (SIGGRAPH ’24). Association for Computing Machinery, New York, NY, USA, Article 2, 2 pages. https://doi.org/10.1145/3641520.3665309
- Huang et al. (2023) Han Huang, Fernanda De La Torre, Cathy Mengying Fang, Andrzej Banburski-Fahey, Judith Amores, and Jaron Lanier. 2023. Real-time animation generation and control on rigged models via large language models. arXiv preprint arXiv:2310.17838 (2023).
- Izhikevich et al. (2015) Eugene M Izhikevich, John H Conway, and Anil Seth. 2015. Game of life. Scholarpedia 10, 6 (2015), 1816.
- Jiang et al. (2023) Albert Qiaochu Jiang, Alexandre Sablayrolles, Arthur Mensch, Chris Bamford, Devendra Singh Chaplot, Diego de Las Casas, Florian Bressand, Gianna Lengyel, Guillaume Lample, Lucile Saulnier, L’elio Renard Lavaud, Marie-Anne Lachaux, Pierre Stock, Teven Le Scao, Thibaut Lavril, Thomas Wang, Timothée Lacroix, and William El Sayed. 2023. Mistral 7B. ArXiv abs/2310.06825 (2023). https://api.semanticscholar.org/CorpusID:263830494
- Karth and Compton (2023) Isaac Karth and Kate Compton. 2023. Conceptual Art Made Real: Why Procedural Content Generation is Impossible. In Proceedings of the 18th International Conference on the Foundations of Digital Games. 1–3.
- Lanzi and Loiacono (2023) Pier Luca Lanzi and Daniele Loiacono. 2023. Chatgpt and other large language models as evolutionary engines for online interactive collaborative game design. In Proceedings of the Genetic and Evolutionary Computation Conference. 1383–1390.
- Liapis et al. (2013) Antonios Liapis, Georgios N Yannakakis, and Julian Togelius. 2013. Sentient sketchbook: computer-assisted game level authoring. (2013).
- LING et al. (2024) LONG LING, XINYI CHEN, RUOYU WEN, TOBY JIA-JUN LI, and RAY LC. 2024. Sketchar: Supporting Character Design and Illustration Prototyping Using Generative AI. (2024).
- Liu et al. (2023) Vivian Liu, Jo Vermeulen, George Fitzmaurice, and Justin Matejka. 2023. 3DALL-E: Integrating text-to-image AI in 3D design workflows. In Proceedings of the 2023 ACM designing interactive systems conference. 1955–1977.
- Lozhkov et al. (2024) Anton Lozhkov, Raymond Li, Loubna Ben Allal, Federico Cassano, Joel Lamy-Poirier, Nouamane Tazi, Ao Tang, Dmytro Pykhtar, Jiawei Liu, Yuxiang Wei, Tianyang Liu, Max Tian, Denis Kocetkov, Arthur Zucker, Younes Belkada, Zijian Wang, Qian Liu, Dmitry Abulkhanov, Indraneil Paul, Zhuang Li, Wen-Ding Li, Megan L. Risdal, Jia Li, Jian Zhu, Terry Yue Zhuo, Evgenii Zheltonozhskii, Nii Osae Osae Dade, W. Yu, Lucas Krauss, Naman Jain, Yixuan Su, Xuanli He, Manan Dey, Edoardo Abati, Yekun Chai, Niklas Muennighoff, Xiangru Tang, Muhtasham Oblokulov, Christopher Akiki, Marc Marone, Chenghao Mou, Mayank Mishra, Alexander Gu, Binyuan Hui, Tri Dao, Armel Randy Zebaze, Olivier Dehaene, Nicolas Patry, Canwen Xu, Julian J. McAuley, Han Hu, Torsten Scholak, Sébastien Paquet, Jennifer Robinson, Carolyn Jane Anderson, Nicolas Chapados, Mostofa Patwary, Nima Tajbakhsh, Yacine Jernite, Carlos Muñoz Ferrandis, Lingming Zhang, Sean Hughes, Thomas Wolf, Arjun Guha, Leandro von Werra, and Harm de Vries. 2024. StarCoder 2 and The Stack v2: The Next Generation. ArXiv abs/2402.19173 (2024). https://api.semanticscholar.org/CorpusID:268063676
- McDonald et al. (2019) Nora McDonald, Sarita Schoenebeck, and Andrea Forte. 2019. Reliability and inter-rater reliability in qualitative research: Norms and guidelines for CSCW and HCI practice. Proceedings of the ACM on human-computer interaction 3, CSCW (2019), 1–23.
- Nasir et al. (2024) Muhammad U Nasir, Steven James, and Julian Togelius. 2024. Word2World: Generating Stories and Worlds through Large Language Models. arXiv preprint arXiv:2405.06686 (2024).
- Nichol et al. (2021) Alex Nichol, Prafulla Dhariwal, Aditya Ramesh, Pranav Shyam, Pamela Mishkin, Bob McGrew, Ilya Sutskever, and Mark Chen. 2021. GLIDE: Towards Photorealistic Image Generation and Editing with Text-Guided Diffusion Models. In International Conference on Machine Learning. https://api.semanticscholar.org/CorpusID:245335086
- O’Neill et al. (2009) Brian O’Neill, Mark O. Riedl, and Michael Nitsche. 2009. Towards intelligent authoring tools for machinima creation. In CHI ’09 Extended Abstracts on Human Factors in Computing Systems (Boston, MA, USA) (CHI EA ’09). Association for Computing Machinery, New York, NY, USA, 4639–4644. https://doi.org/10.1145/1520340.1520713
- OpenAI (2024a) OpenAI. 2024a. ChatGPT: Language Model. https://chat.openai.com/. Accessed: 2024-09-02.
- OpenAI (2024b) OpenAI. 2024b. Hello GPT-4o. https://openai.com/index/hello-gpt-4o/ Accessed: 2024-09-11.
- Panchanadikar et al. (2024) Ruchi Panchanadikar, Guo Freeman, Lingyuan Li, Kelsea Schulenberg, and Yang Hu. 2024. ”A New Golden Era” or ”Slap Comps”: How Non-Profit Driven Indie Game Developers Perceive the Emerging Role of Generative AI in Game Development. In Extended Abstracts of the 2024 CHI Conference on Human Factors in Computing Systems (CHI EA ’24). Association for Computing Machinery, New York, NY, USA, Article 1, 7 pages. https://doi.org/10.1145/3613905.3650845
- Pugh et al. (2016) Justin K Pugh, Lisa B Soros, and Kenneth O Stanley. 2016. Quality diversity: A new frontier for evolutionary computation. Frontiers in Robotics and AI 3 (2016), 202845.
- Radford et al. (2021) Alec Radford, Jong Wook Kim, Chris Hallacy, Aditya Ramesh, Gabriel Goh, Sandhini Agarwal, Girish Sastry, Amanda Askell, Pamela Mishkin, Jack Clark, et al. 2021. Learning transferable visual models from natural language supervision. In International conference on machine learning. PMLR, 8748–8763.
- Ramesh et al. (2022) Aditya Ramesh, Prafulla Dhariwal, Alex Nichol, Casey Chu, and Mark Chen. 2022. Hierarchical Text-Conditional Image Generation with CLIP Latents. ArXiv abs/2204.06125 (2022). https://api.semanticscholar.org/CorpusID:248097655
- Richardson et al. (2023) Elad Richardson, Gal Metzer, Yuval Alaluf, Raja Giryes, and Daniel Cohen-Or. 2023. Texture: Text-guided texturing of 3d shapes. In ACM SIGGRAPH 2023 conference proceedings. 1–11.
- Rombach et al. (2022) Robin Rombach, Andreas Blattmann, Dominik Lorenz, Patrick Esser, and Björn Ommer. 2022. High-resolution image synthesis with latent diffusion models. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 10684–10695.
- Secretan et al. (2011) Jimmy Secretan, Nicholas Beato, David B D’Ambrosio, Adelein Rodriguez, Adam Campbell, Jeremiah T Folsom-Kovarik, and Kenneth O Stanley. 2011. Picbreeder: A case study in collaborative evolutionary exploration of design space. Evolutionary computation 19, 3 (2011), 373–403.
- SideFX ([n. d.]) SideFX. [n. d.]. Houdini software. https://www.sidefx.com/products/houdini/
- Stanley (2007) Kenneth O Stanley. 2007. Compositional pattern producing networks: A novel abstraction of development. Genetic programming and evolvable machines 8 (2007), 131–162.
- Stanley and Lehman (2015) Kenneth O Stanley and Joel Lehman. 2015. Why greatness cannot be planned: The myth of the objective. Springer.
- Sudhakaran et al. (2024) Shyam Sudhakaran, Miguel González-Duque, Matthias Freiberger, Claire Glanois, Elias Najarro, and Sebastian Risi. 2024. Mariogpt: Open-ended text2level generation through large language models. Advances in Neural Information Processing Systems 36 (2024).
- Sun et al. (2024) Qirui Sun, Qiaoyang Luo, Yunyi Ni, and Haipeng Mi. 2024. Text2AC: A Framework for Game-Ready 2D Agent Character(AC) Generation from Natural Language. In Extended Abstracts of the 2024 CHI Conference on Human Factors in Computing Systems (CHI EA ’24). Association for Computing Machinery, New York, NY, USA, Article 312, 7 pages. https://doi.org/10.1145/3613905.3651049
- Tochilkin et al. (2024) Dmitry Tochilkin, David Pankratz, Zexiang Liu, Zixuan Huang, Adam Letts, Yangguang Li, Ding Liang, Christian Laforte, Varun Jampani, and Yan-Pei Cao. 2024. Triposr: Fast 3d object reconstruction from a single image. arXiv preprint arXiv:2403.02151 (2024).
- Todd et al. (2023) Graham Todd, Sam Earle, Muhammad Umair Nasir, Michael Cerny Green, and Julian Togelius. 2023. Level generation through large language models. In Proceedings of the 18th International Conference on the Foundations of Digital Games. 1–8.
- Torre et al. (2023) Fernanda De La Torre, Cathy Mengying Fang, Han Huang, Andrzej Banburski-Fahey, Judith Amores Fernandez, and Jaron Lanier. 2023. LLMR: Real-time Prompting of Interactive Worlds using Large Language Models. Proceedings of the CHI Conference on Human Factors in Computing Systems (2023). https://api.semanticscholar.org/CorpusID:262084413
- Touvron et al. (2023) Hugo Touvron, Louis Martin, Kevin R. Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, Daniel M. Bikel, Lukas Blecher, Cristian Cantón Ferrer, Moya Chen, Guillem Cucurull, David Esiobu, Jude Fernandes, Jeremy Fu, Wenyin Fu, Brian Fuller, Cynthia Gao, Vedanuj Goswami, Naman Goyal, Anthony S. Hartshorn, Saghar Hosseini, Rui Hou, Hakan Inan, Marcin Kardas, Viktor Kerkez, Madian Khabsa, Isabel M. Kloumann, A. V. Korenev, Punit Singh Koura, Marie-Anne Lachaux, Thibaut Lavril, Jenya Lee, Diana Liskovich, Yinghai Lu, Yuning Mao, Xavier Martinet, Todor Mihaylov, Pushkar Mishra, Igor Molybog, Yixin Nie, Andrew Poulton, Jeremy Reizenstein, Rashi Rungta, Kalyan Saladi, Alan Schelten, Ruan Silva, Eric Michael Smith, R. Subramanian, Xia Tan, Binh Tang, Ross Taylor, Adina Williams, Jian Xiang Kuan, Puxin Xu, Zhengxu Yan, Iliyan Zarov, Yuchen Zhang, Angela Fan, Melanie Kambadur, Sharan Narang, Aurelien Rodriguez, Robert Stojnic, Sergey Edunov, and Thomas Scialom. 2023. Llama 2: Open Foundation and Fine-Tuned Chat Models. ArXiv abs/2307.09288 (2023). https://api.semanticscholar.org/CorpusID:259950998
- Valevski et al. (2024) Dani Valevski, Yaniv Leviathan, Moab Arar, and Shlomi Fruchter. 2024. Diffusion Models Are Real-Time Game Engines. arXiv:2408.14837 [cs.LG] https://arxiv.org/abs/2408.14837
- Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed Chi, Quoc V Le, Denny Zhou, et al. 2022. Chain-of-thought prompting elicits reasoning in large language models. Advances in neural information processing systems 35 (2022), 24824–24837.
- Wu et al. (2020) Zonghan Wu, Shirui Pan, Fengwen Chen, Guodong Long, Chengqi Zhang, and S Yu Philip. 2020. A comprehensive survey on graph neural networks. IEEE transactions on neural networks and learning systems 32, 1 (2020), 4–24.
- Yang et al. (2024) Daijin Yang, Erica Kleinman, and Casper Harteveld. 2024. GPT for Games: A Scoping Review (2020-2023). arXiv preprint arXiv:2404.17794 (2024).
- Zhao et al. (2024) CodeGemma Team Heri Zhao, Jeffrey Hui, Joshua Howland, Nam Nguyen, Siqi Zuo, Andrea Hu, Christopher A. Choquette-Choo, Jingyue Shen, Joe Kelley, Kshi tij Bansal, Luke Vilnis, Mateo Wirth, Paul Michel, Peter Choy, Pratik Joshi, Ravin Kumar, Sarmad Hashmi, Shubham Agrawal, Zhitao Gong, Jane Fine, Tris Brian Warkentin, Ale Jakse Hartman, Bin Ni, Kathy Korevec, Kelly Schaefer, and Scott Huffman. 2024. CodeGemma: Open Code Models Based on Gemma. ArXiv abs/2406.11409 (2024). https://api.semanticscholar.org/CorpusID:270560319
Appendix A Prompts
Here we provide the prompts used by some of the modules in DreamGarden.
A.1. Planner
Here is the python code used to generate system prompts for the planner module (both when transforming the initial user seed prompt to a broad plan, and when recursively breaking down the plan into more fine-grained steps). Note that the maximum branching factor and depth are specified in the config ‘cfg‘. In our experiments, ‘cfg.test_disclaimer‘ is always ‘True‘.
Here is the code used to generate prompts based on user input and the existing plan tree.
A.2. Task generator
The following code produces prompts used by the task generator, which is responsible for transforming leaf nodes in the tree generated by the planner into concrete implementation tasks (i.e. into prompts for specific implementation submodules). Note that this code makes use of the ‘available_submodules‘ dictionary defined above.
A.3. Code generator
The following code produces prompts for generating/repairing/extending code, including in cases where the code should involve a procedural mesh.
The following code produces prompts for evaluating the compilation logs produced when compiling generated code, and providinv feedback to the code generator in case of compilation failure.
The following code produces prompts for evaluating the Unreal Engine crash logs—in case a run-time crash occurs when running generated code—and providing feedback to the code generator.
The following code produces prompts for evaluating visual output and runtime logs in case generated code compiles and runs without error.
Appendix B User Study
B.1. Intake survey
What is your level of familiarity with Generative AI systems like ChatGPT or DALL-E Image Generator?
Can you briefly describe your experience using Generative AI systems? (N/A if not applicable)
What is your level of familiarity with game design and/or game engines?
Can you briefly describe your experience with game design and/or game engines?
(N/A if not applicable) What is your level of familiarity with Unreal Engine?
Can you briefly describe your experience with Unreal Engine? (N/A if not applicable)
Here, questions asking for a user’s level of familiarity have multiple choice answers with the following options:
Very unfamiliar
Unfamiliar
Slightly unfamiliar
Slightly familiar
Familiar
Very familiar
Other questions allowed for short responses.
B.2. Exit survey
In what scenarios would you find it valuable to use a tool like LucidDev333Note that during the user study, the system had the working name “LucidDev”, which was later changed to “DreamGarden”.?
Are there any situations where you would prefer not to use a tool like LucidDev? If so, why?
What aspects of the system, if any, did you find appealing/enjoyable? What aspects did you find unappealing/unenjoyable?
Reflecting on LucidDev in its current state, imagine a more fully-fledged version of this system. In a perfect world, what would this system look like? What features would the system have that it currently lacks?
Brainstorm positive and negative implications of your vision. Describe them here.
On a scale from 1-5 (5 being most enjoyable) how enjoyable did you find using the tool?
On a scale from 1-5 (5 being most frustrating) how frustrating did you find using the tool?
All questions in the exit survey allowed for short responses from users.